[figure: 9.7 cm]

MAGNETIC ORDERING OF RAgSb_{2} COMPOUNDS

G. André^{a}, F. Bourée^{a}, M. Kolenda^{b}, B. Leśniewska^{c}, A. Oleś^{c} and A. Szytuła^{b} a Laboratoire Léon Brillouin (CEA-CNRS) CEN-Saclay, 91191 Gif-Sur-Yvette Cedex, France

b Institute of Physics, Jagiellonian University, Reymonta 4, 30-059 Kraków, Poland c Department of Physics and Nuclear Techniques, Academy of Mining and Metallurgy,

Reymonta 19, 30-059 Kraków, Poland

Results of neutron diffraction measurements for RAgSb_{2} (R = Ce, Pr, Nd, Tb-Tm) compounds are presented. All the compounds crystallize in the tetragonal ZrCuSi_{2}-type of crystal structure and except for CeAgSb_{2} they are antiferromagnets with the Néel temperature 11.0 K for R=Tb and 2.0 K for R=Tm. The CeAgSb_{2} below T_{C}=9.5 K is a ferromagnet with the magnetic moment equal to 0.33 μ_{B} and parallel to the c-axis. Below T_{N} PrAgSb_{2} has a collinear antiferromagnetic structure of AF II type with the magnetic unit cell (a,a,2c). In NdAgSb_{2} a sinemodulated magnetic ordering described by the wave vector k=(0.5435,0,0) is observed. The compounds with R=Tb,Dy and Er below T_{N} have a collinear structure with a magnetic unit cell doubled along the a-axis respective to the crystal one. The magnetic structure of the Ho-compounds at T =1.4 K is similar to that observed in TbAgSb_{2} and DyAgSb_{2}. When the temperature increases a commensurate collinear and an incommensurate sine modulated structures coexist. The two phases are observed at T =1.4 K also in TmAgSb_{2}. An increase of the temperature causes that the commensurate phase disappears. The magnetic moment is parallel to the b-axis for R=Tb and Dy compounds and parallel to the c-axis for R=Er and Tm which correlates with the signs B-parameter. In HoAgSb_{2} magnetic moment forms angle ϕ_{b}=13.5^{0} with the b-axis. An increase of temperature changes the angle ϕ_{b}.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds Presentation mode : oral Corresponding author : Prof. dr hab. Andrzej Szytula Address for correspondence : Instytut Fizyki UJ 30-059 Krakow, Reymonta 4 tel (12) 632 48 88 / 5546 fax: (12) 633 70 86 Email address : szytula@if.uj.edu.pl